OpenStudy (anonymous):

evaluate integral sin^3x cos^2xdx

OpenStudy (amistre64):

change sin^3 to sin sin^2; sin^2 = 1-cos^2; and fit that to the cos^2 already there?

OpenStudy (anonymous):

yup

OpenStudy (amistre64):

\[\int sin\ sin^2\ cos^2\] \[\int sin\ (1-cos^2)\ cos^2\] \[\int sin\ (cos^2-cos^4)\] \[\int (cos^2sin-cos^4sin)\] perhaps?

OpenStudy (anonymous):

yes but is there something that u= cosx and du= -sinx?

OpenStudy (amistre64):

\[\int cos^2sin-\int cos^4sin\] \[ -\frac{cos^3}{3}+ \frac{cos^5}{5}\]

OpenStudy (anonymous):

let cox = z then dz = -sinx dx now solve

OpenStudy (amistre64):

sure, but only after you split it into workable parts

OpenStudy (anonymous):

ok thanks!! big help : )

OpenStudy (amistre64):

i think the wolfram disagrees with my assessment tho

OpenStudy (amistre64):

so i could have made a mistake somewhere

OpenStudy (anonymous):

\(\int\limits_{}^{}\sin^3{x} \cos^2{x}dx=\int\limits_{}^{}\sin{x}(\cos^2x-\cos^4{x})dx\). Use a substitution \(u=\cos{x} \implies du=-\sin{x}dx\). So the integral becomes \(\int\limits_{}^{}(u^4-u^2)du={u^5 \over 5}-{u^3 \over 3}+c\). Now, just substitute back \(u=\cos{x}\).

OpenStudy (amistre64):

no its good, they just take it another step farther

OpenStudy (anonymous):

(1-cos^2 x)cos^2 x (-dz) => -(1-z^2)(z^2)dz => (-z^2 + z^4)dz => -z^3/3 + z^5/5 => - cos^3/3 + cos^5/5 +c

OpenStudy (anonymous):

I think I am too slow! -.-

OpenStudy (amistre64):

the wolf give ...which i do not see as simplier by any means 1/30 cos^3(x) (3 cos(2 x)-7)+c

OpenStudy (amistre64):

its not that there is something to do with a u-sub; there is nothing magical about the usub other than it tends to clean up the clutter